# Transolver is a Linear Transformer: Revisiting Physics-Attention through the Lens of Linear Attention

### v2

@@ Introduction
Solving Partial Differential Equations is a fundamental task in many fields of science and engineering. Due to their complexity, these equations often require discretization into meshes and solving with numerical methods (piomelli1999large; alfonsi2009reynolds; lee2015direct), which are computationally expensive and time-consuming. Recent advances in deep learning provide new approaches for PDE solving (chen2022flowdnn; liu2024lkflownet; huang2025partial).
The essence of PDE solving is to find the solution function corresponding to the PDE and its boundary or initial conditions. This can be formulated as a supervised learning problem, where neural networks learn the mappings from boundary or initial functions to solution functions. Neural Operator (luLearningNonlinearOperators2021; kovachki2023neural) was proposed to learn the mappings between those two function spaces while guaranteeing both discretization-invariance and universal approximation.
@@
The integral kernel operator of Neural Operator (kovachki2023neural) is a non-local operation that passes messages between discrete grid points in a discretization-invariant manner, which is analogous to the transformer’s self-attention (vaswani2017attention). Therefore, a series of works have introduced transformers into PDE solving and Neural Operator construction. However, when considering discrete grid points as tokens, the quadratic complexity of self-attention seriously limits the scale of PDE problems. Several strategies are employed to optimize the complexity of self-attention. Some methods (li2023transformer; li2023scalable; haoGNOTGeneralNeural2023) introduce linear attention  (DBLP:conf/icml/KatharopoulosV020) to decrease the quadratic complexity to linear. Other methods (wangLatentNeuralOperator2024; serranoAROMAPreservingSpatial2024; alkinUniversalPhysicsTransformers2024) follow the Set Transformer (lee2019set), projecting grid points into a latent space to reduce their length and performing attention in that space for PDE solving. All of the above works aim to construct low-rank patterns to achieve complexity reduction. Transolver (wu2024transolver), as a representative model, proposes a Physics-Attention module which contains a slicing operation to project the grid points into several slices and perform slice attention, and a deslicing operation to project the slices back to grid points. However, the design of this module is still intuitive, and its rationality lacks in-depth analysis.
@@
Neural Operators are a class of models that aim to solve PDE by learning a mapping from boundary or initial functions to solution functions. They achieve mappings between function spaces by ensuring universal approximation and discretization-invariance (kovachki2023neural). DeepONet (luLearningNonlinearOperators2021) uses a two-branch architecture to learn the mapping between functions. Graph Neural Operator (GNO) (liNeuralOperatorGraph2020) implements the integral kernel operator using message passing in graph neural networks. Fourier Neural Operator (FNO) (liFourierNeuralOperator2021) performs the integral kernel operator by mapping input functions to the frequency domain via Fourier transforms. GEO-FNO (DBLP:journals/jmlr/LiHLA23) extends FNO to unstructured meshes by mapping irregular domains onto structured grids. U-FNO (DBLP:journals/corr/abs-2109-03697) enhances the multi-scale feature extraction capability of FNO by incorporating a U-Net architecture.
@@
Transformer (vaswani2017attention) has been shown to be a special case of integral kernel operators (kovachki2023neural). However, its quadratic computational complexity makes it difficult to apply to large-scale PDE problems efficiently. Some methods, such as OFormer (li2023transformer), FactFormer (li2023scalable), and GNOT (haoGNOTGeneralNeural2023), adopt linear attention to reduce the computational cost while maintaining the modeling capacity. Another line of methods aims to reduce complexity by compressing the length of the token sequence. Methods such as LNO (wangLatentNeuralOperator2024), AROMAP (serranoAROMAPreservingSpatial2024), UPT (alkinUniversalPhysicsTransformers2024), and AeroGTO (liuAeroGTOEfficientGraphTransformer) reduce computational complexity by introducing a learnable set of latent tokens to project discrete grid points into a latent space using cross-attention, followed by self-attention within the latent space. LSM (wuSolvingHighDimensionalPDEs2023) also employs the same compression strategy to reduce sequence length, and replaces self-attention with spectral transformation for compressed tokens. Transolver (wu2024transolver) introduces Physics-Attention, which uses a learnable slice-weight matrix to assign discrete grid points to physical slices, grouping similar grid points together. Self-attention is then applied across slices to exchange information and uncover deeper physical relationships within the complex geometry. The slice-weight matrix is used to decode the slices back to grid points. Transolver++ (DBLP:journals/corr/abs-2502-02414) proposes local adaptation for better slice distinction.
@@
The Softmax operation on the attention matrix is the key factor contributing to the quadratic complexity of attention. To address this issue, linear attention uses kernel functions to approximate the Softmax function, allowing it to change the computation order of $\bm{Q}$, $\bm{K}$, and $\bm{V}$ in full attention, thereby reducing the computational complexity from quadratic to linear. The Linear Transformer (DBLP:conf/icml/KatharopoulosV020) replaces the softmax operation with a kernel function, achieving performance comparable to standard attention. Both Performer (DBLP:journals/corr/abs-2006-03555) and RFA  (DBLP:conf/iclr/Peng0Y0SK21) employ random feature methods to approximate the softmax function, making the approximation mathematically equivalent to full attention. TSSA (DBLP:conf/iclr/WuDLPZWY0H25) derives a new linear attention by optimizing a variational form of the Maximum Coding Rate Reduction objective. RALA (fan2025breaking) introduces a rank-augmented method to enhance the performance.
@@
Transolver (wu2024transolver) proposes a Transformer-based Neural Operator to capture the physical properties of PDE solutions through the Physics-Attention. In this section, we briefly revisit its core components. The input to the Physics-Attention is a feature matrix $\bm{H}_{N}=\{\bm{h}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{h}}$, obtained from $\bm{A}_{N}$ and $\bm{X}_{N}$, where $N$ is the number of spatial grid points and $d_{h}$ is the feature dimension. First, the input feature matrix $\bm{H}_{N}$ is passed through a linear layer followed by a softmax function to produce a slice-weight matrix $\bm{W}_{N}=\{\bm{w}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times M}$, where $M$ is a hyperparameter denoting the number of slices. A channel-wise linear layer is subsequently applied to $\bm{H}_{N}$ to obtain $\bm{V}_{N}=\{\bm{v}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{h}}$. Using this slice-weight matrix, Physics-Attention computes a weighted average of $\bm{V}_{N}$ to obtain a slice feature matrix $\bm{S}_{M}=\{\bm{s}_{j}\}_{j=1}^{M}\in\mathbb{R}^{M\times d_{h}}$. This operation can be interpreted as projecting the original $N$ grid points onto $M$ physical slices, effectively compressing the sequence length from $N$ to $M$. The computation is formally given by:
@@
The $\text{Linear}^{(Q/K)}(\cdot)$ and $\text{Linear}^{(V)}(\cdot)$ are single-layer MLPs with output dimensions of $M$ and $d_{h}$, respectively. A self-attention on the slices $\bm{S}_{M}$ is then performed to produce an updated slices $\bm{S^{\prime}}_{M}=\{\bm{s^{\prime}}_{j}\}_{j=1}^{M}\in\mathbb{R}^{M\times d_{h}}$. Finally, using the slice-weight matrix $\bm{W}_{N}$, the slices $\bm{S^{\prime}}_{M}$ are decoded back to the original sequence length, yielding the output $\bm{H^{\prime}}_{N}=\{\bm{h^{\prime}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{h}}$. The operations can be summarized as follows:
@@
As shown in Eq. (6), linear attention aims to approximate the softmax function by designing different kernel functions $\varphi(\cdot)$ and $\psi(\cdot)$, allowing it to change the computation order of $\bm{Q}$, $\bm{K}$, and $\bm{V}$. In OFormer (li2023transformer) and FactFormer (li2023scalable), the kernel functions correspond to vector normalization and the identity function, respectively. The choice of kernel functions directly affects the performance of linear attention.
@@
Some improvements of linear attention insert additional intermediate operations between $\psi^{\top}(\bm{K})\bm{V}$ and $\varphi(\bm{Q})$, i.e., $\varphi(\bm{Q})\circ\mathcal{G}\circ\psi^{\top}(\bm{K})\bm{V}$. The operation $\mathcal{G}$ can be rank augmentation (fan2025breaking), forget gate (yang2023gated), etc.
@@
In this section, we present a more generalized analysis of the Physics-Attention. Based on this, we redesign it through two key steps—generalization and simplification—to derive a more flexible architecture called the Linear Attention Neural Operator, or LinearNO. Figure 2(b) shows a comparison between the proposed LinearNO and the original Physics-Attention.
@@
DBLP:journals/corr/abs-2502-02414 (DBLP:journals/corr/abs-2502-02414) observed that Physics-Attention tends to produce uniform slice weights, thus generating less distinguishable slices in some cases, which damages the model’s performance, as shown in Figure 3. We hypothesize that this may be attributed to the intended asymmetry between $\varphi(\bm{Q})$ and $\psi(\bm{K})$ in linear attention; enforcing the shared learnable layer between $\varphi(\bm{Q})$ and $\psi(\bm{K})$ could blur their roles and result in overly averaged representations. Based on this hypothesis, we relaxed the weight-sharing constraint in the generalization step, resulting in asymmetric $\varphi(\bm{Q})$ and $\psi(\bm{K})$. Specifically, we allow $\psi(\bm{K})$ to be learned independently of $\varphi(\bm{Q})$. Following Physics-Attention, $\varphi(\bm{Q})$ and $\psi(\bm{K})$ are normalized along the $M$ and $N$ dimensions, respectively. This ensures that each row of attention matrix $\varphi(\bm{Q})\psi^{\top}(\bm{K})$ is normalized, which is consistent with full attention (shen2021efficient). Formally, we define $\varphi(\bm{Q})$ and $\psi(\bm{K})$ as follows:
@@
FNO (liFourierNeuralOperator2021) / / / 0.1556 0.0108 /
U-FNO (DBLP:journals/corr/abs-2109-03697) 0.0269 0.0056 0.0039 0.2231 0.0183 0.0239
GEO-FNO (DBLP:journals/jmlr/LiHLA23) 0.0138 0.0067 0.0074 0.1556 0.0108 0.0229
F-FNO (DBLP:conf/iclr/TranMXO23) 0.0078 0.0070 0.0047 0.2322 0.0077 0.0263
Galerkin (DBLP:conf/nips/Cao21) 0.0118 0.0098 0.0120 0.1401 0.0084 0.0240
OFormer (li2023transformer) 0.0183 0.0168 0.0017 0.1705 0.0124 0.0183
GNOT (haoGNOTGeneralNeural2023) 0.0076 0.0047 0.0336 0.1380 0.0105 0.0086
FactFormer (li2023scalable) 0.0071 0.0060 0.0312 0.1214 0.0109 /
ONO (DBLP:conf/icml/XiaoHLD024) 0.0061 0.0052 0.0048 0.1195 0.0076 0.0118
LSM (wuSolvingHighDimensionalPDEs2023) 0.0059 0.0050 0.0025 0.1535 0.0065 0.0218
LNO (wangLatentNeuralOperator2024) 0.0053 0.0031 0.0028 0.0830 0.0063 0.0066
Transolver (wu2024transolver) 0.0053 0.0030 0.0013 0.0882 0.0055 0.0065
Transolver++ (DBLP:journals/corr/abs-2502-02414) 0.0051 0.0027 0.0014 0.1010 0.0056 0.0064
@@
GraphSage (DBLP:conf/nips/HamiltonYL17) 0.0087 0.0184 0.1476 0.9964 0.0461 0.1050 0.0270 0.9695
PointNet (DBLP:conf/cvpr/QiSMG17) 0.0253 0.0996 0.1973 0.9919 0.0494 0.1104 0.0298 0.9583
GraphUNet (DBLP:conf/icml/GaoJ19) 0.0076 0.0144 0.1677 0.9949 0.0471 0.1102 0.0226 0.9725
MeshGraphNet (DBLP:conf/iclr/PfaffFSB21) 0.0214 0.0387 0.2252 0.9945 0.0354 0.0781 0.0168 0.9840
GNO (liNeuralOperatorGraph2020) 0.0269 0.0405 0.2016 0.9938 0.0383 0.0815 0.0172 0.9834
GEO-FNO (DBLP:journals/jmlr/LiHLA23) 0.0361 0.0301 0.6161 0.9257 0.1670 0.2378 0.0664 0.8280
GALERKIN (DBLP:conf/nips/Cao21) 0.0074 0.0159 0.2336 0.9951 0.0339 0.0878 0.0179 0.9764
GNOT (haoGNOTGeneralNeural2023) 0.0049 0.0152 0.1992 0.9942 0.0329 0.0798 0.0178 0.9833
GINO (DBLP:conf/nips/LiKCLKONS0AA23) 0.0297 0.0482 0.1821 0.9958 0.0386 0.0810 0.0184 0.9826
LNO (wangLatentNeuralOperator2024) 0.0214 0.0268 0.1480 0.9744 0.0269 0.0870 0.0174 0.9781
Transolver (wu2024transolver) 0.0023 0.0085 0.1230 0.9978 0.0221 0.0785 0.0117 0.9914
Transolver++ (DBLP:journals/corr/abs-2502-02414) 0.0068 0.0159 0.1880 0.9910 0.0226 0.0800 0.0132 0.9914
@@
Datasets We adopt six benchmark datasets, including Airfoil, Pipe, Plasticity, NS, Darcy and Elasticity, which are classical problems in fluid mechanics and solid mechanics. These datasets were introduced by FNO (liFourierNeuralOperator2021) and Geo-FNO (DBLP:journals/jmlr/LiHLA23), and are now widely used as standard benchmarks for evaluating Neural Operators. In addition, we use two industrial-level datasets: AirfRANS (bonnet2022airfrans) and ShapeNet-Car (DBLP:journals/tog/ShapeCar) to assess the model’s performance in aerodynamic shape design. More detailed information about all datasets can be found in Appendix B.
Baselines We compare our model against several representative baselines, including classical Neural Operator models: FNO (liFourierNeuralOperator2021), Geo-FNO (DBLP:journals/jmlr/LiHLA23), F-FNO (DBLP:conf/iclr/TranMXO23) and U-FNO (DBLP:journals/corr/abs-2109-03697), Transformer-based models: Galerkin (DBLP:conf/nips/Cao21), OFormer (li2023transformer), GNOT (haoGNOTGeneralNeural2023), FactFormer (li2023scalable), ONO (DBLP:conf/icml/XiaoHLD024), LSM (wuSolvingHighDimensionalPDEs2023), LNO (wangLatentNeuralOperator2024), Transolver (wu2024transolver) and Transolver++ (DBLP:journals/corr/abs-2502-02414), classical geometric deep models: PointNet (DBLP:conf/cvpr/QiSMG17), GraphSage (DBLP:conf/nips/HamiltonYL17) and GraphUNet (DBLP:conf/icml/GaoJ19), MeshGraphNet (DBLP:conf/iclr/PfaffFSB21).
Setup To ensure fair comparisons, we follow the experimental settings of Transolver (wu2024transolver). We report the relative L2 error as the evaluation metric, except the Spearman’s correlation coefficient $\rho$. For full experimental details and hyperparameter configurations, please refer to Appendix B.
Figure 4: The left plot shows the visualization of the final-layer $\varphi(\bm{Q})$ generated by LinearNO and Transolver. The middle plot visualizes the error distributions of LinearNO and Transolver. The right plot illustrates two metrics: the average rank of the attention matrix $\varphi(\bm{Q})\psi^{\top}(\bm{K})$ per head at each layer when Slice $M=64$, and the prediction errors of both models on the Airfoil dataset under different slice numbers.
@@
Table 1 presents the relative L2 errors of LinearNO and SOTA models across six PDE benchmarks. LinearNO achieves the best performance on all tasks. Specifically, LinearNO achieves over a 10% relative improvement on the Pipe, Plasticity, NS, Darcy and Elasticity task.
@@
We compare the number of parameters and computational cost between LinearNO and Transolver on six standard PDE tasks, as shown in Table 3. It can be observed that LinearNO has significantly fewer parameters and lower computational cost than Transolver. On average, our method reduces the number of parameters by 40.0% and the FLOPs by 36.2%. This demonstrates that LinearNO is more computationally efficient and lightweight, making it more suitable for deployment in resource-constrained environments.
@@
Slice Analysis.
The rank of attention matrix $\varphi(\bm{Q})\psi^{\top}(\bm{K})$ corresponds to the number of physical states effectively captured by the model, which influences its performance (DBLP:journals/corr/abs-2502-02414). This effective rank is bounded by the number of slices $M$. We hypothesize that the symmetric design of $\varphi(\bm{Q})$ and $\psi(\bm{K})$ in Transolver limits the effective utilization of slices. In this section, we present a quantitative analysis to support this claim.
We estimate the layer-wise rank of the attention matrix using Singular Value Decomposition (Golub1965SVD), and the results are shown in Figure 4. The findings indicate that the asymmetric design of $\varphi(\bm{Q})$ and $\psi(\bm{K})$ facilitates more efficient utilization of slices. Furthermore, we visualize the distribution of the final-layer $\varphi(\bm{Q})$ matrix on the Airfoil dataset. Compared to Transolver, LinearNO exhibits more diverse physical slices. In addition, LinearNO consistently outperforms Transolver across different numbers of slices on the Airfoil dataset.
@@
We design ablation experiments to evaluate the impact of the generalization and simplification steps on model performance. Specifically, we set $\psi^{\top}(\bm{K})=\text{Softmax}(\varphi^{\top}(\bm{Q}))$ to disable the generalization step, and we add a slice attention to disable the simplification step. We compare these variants with LinearNO, as shown in Table 4. By comparing No.1 vs. No.2 and No.3 vs. No.4, we observe that the generalization step consistently improves model performance in most cases, confirming the benefit of using asymmetric projections. Furthermore, the comparison between No.1 and No.3 shows that, under symmetric settings, slice attention does not consistently improve performance, which is also observed in Transolver. In contrast, comparing No.2 and No.4 reveals that applying slice attention under asymmetric projections consistently hurts the performance. Therefore, the simplification step, which disables slice attention, is also justified.

### v3

@@ Introduction
Solving Partial Differential Equations is a fundamental task in many fields of science and engineering. Due to their complexity, these equations often require discretization into meshes and solving with numerical methods (Piomelli 1999; Alfonsi 2009; Lee and Moser 2015), which are computationally expensive and time-consuming. Recent advances in deep learning provide new approaches for PDE solving (Chen et al. 2022; Liu et al. 2024; Huang et al. 2025).
The essence of PDE solving is to find the solution function corresponding to the PDE and its boundary or initial conditions. This can be formulated as a supervised learning problem, where neural networks learn the mappings from boundary or initial functions to solution functions. Neural Operator (Lu et al. 2021; Kovachki et al. 2023) was proposed to learn the mappings between those two function spaces while guaranteeing both discretization-invariance and universal approximation.
@@
The integral kernel operator of Neural Operator (Kovachki et al. 2023) is a non-local operation that passes messages between discrete grid points in a discretization-invariant manner, which is analogous to the transformer’s self-attention (Vaswani et al. 2017). Therefore, a series of works have introduced transformers into PDE solving and Neural Operator construction. However, when considering discrete grid points as tokens, the quadratic complexity of self-attention seriously limits the scale of PDE problems. Several strategies are employed to optimize the complexity of self-attention. Some methods (Li et al. 2023a, b; Hao et al. 2023) introduce linear attention  (Katharopoulos et al. 2020) to decrease the quadratic complexity to linear. Other methods (Wang and Wang 2024; Serrano et al. 2024; Alkin et al. 2024) follow the Set Transformer (Lee et al. 2019), projecting grid points into a latent space to reduce their length and performing attention in that space for PDE solving. All of the above works aim to construct low-rank patterns to achieve complexity reduction. Transolver (Wu et al. 2024), as a representative model, proposes a Physics-Attention module which contains a slicing operation to project the grid points into several slices and perform slice attention, and a deslicing operation to project the slices back to grid points. However, the design of this module is still intuitive, and lacks in-depth analysis.
@@
Neural Operators are a class of models that aim to solve PDE by learning a mapping from boundary or initial functions to solution functions. They achieve mappings between function spaces by ensuring universal approximation and discretization-invariance (Kovachki et al. 2023). DeepONet (Lu et al. 2021) uses a two-branch architecture to learn the mapping between functions. Graph Neural Operator (GNO) (Li et al. 2020) implements the integral kernel operator using message passing in graph neural networks. Fourier Neural Operator (FNO) (Li et al. 2021) performs the integral kernel operator by mapping input functions to the frequency domain via Fourier transforms. GEO-FNO (Li et al. 2023c) extends FNO to unstructured meshes by mapping irregular domains onto structured grids. U-FNO (Wen et al. 2022) enhances the multi-scale feature extraction capability of FNO by incorporating a U-Net architecture.
@@
Transformer (Vaswani et al. 2017) has been shown to be a special case of integral kernel operators (Kovachki et al. 2023). However, its quadratic computational complexity makes it difficult to apply to large-scale PDE problems efficiently. Some methods, such as OFormer (Li et al. 2023a), FactFormer (Li et al. 2023b), and GNOT (Hao et al. 2023), adopt linear attention to reduce the computational cost while maintaining the modeling capacity. Another line of methods aims to reduce complexity by compressing the length of the token sequence. Methods such as LNO (Wang and Wang 2024), AROMAP (Serrano et al. 2024), UPT (Alkin et al. 2024), and AeroGTO (Liu et al. 2025) reduce computational complexity by introducing a learnable set of latent tokens to project discrete grid points into a latent space using cross-attention, followed by self-attention within the latent space. LSM (Wu et al. 2023) also employs the same compression strategy to reduce sequence length, and replaces self-attention with spectral transformation for compressed tokens. Transolver (Wu et al. 2024) introduces Physics-Attention, which uses a learnable slice-weight matrix to assign discrete grid points to physical slices, grouping similar grid points together. Self-attention is then applied across slices to exchange information and uncover deeper physical relationships within the complex geometry. The slice-weight matrix is used to decode the slices back to grid points. Transolver++ (Luo et al. 2025) proposes local adaptation for better slice distinction.
@@
The Softmax operation on the attention matrix is the key factor contributing to the quadratic complexity of attention. To address this issue, linear attention uses kernel functions to approximate the Softmax function, allowing it to change the computation order of $\bm{Q}$, $\bm{K}$, and $\bm{V}$ in full attention, thereby reducing the computational complexity from quadratic to linear. The Linear Transformer (Katharopoulos et al. 2020) replaces the softmax operation with a kernel function, achieving performance comparable to standard attention. Both Performer (Choromanski et al. 2020) and RFA  (Peng et al. 2021) employ random feature methods to approximate the softmax function, making the approximation mathematically equivalent to full attention. TSSA (Wu et al. 2025) derives a new linear attention by optimizing a variational form of the Maximum Coding Rate Reduction objective. RALA (Fan et al. 2025) introduces a rank-augmented method to enhance the performance.
@@
Transolver (Wu et al. 2024) proposes a Transformer-based Neural Operator to capture the physical properties of PDE solutions through the Physics-Attention. In this section, we briefly revisit its core components. The input to the Physics-Attention is a feature matrix $\bm{H}_{N}=\{\bm{h}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{h}}$, obtained from $\bm{A}_{N}$ and $\bm{X}_{N}$, where $N$ is the number of spatial grid points and $d_{h}$ is the feature dimension. First, the input feature matrix $\bm{H}_{N}$ is passed through a linear layer followed by a softmax function to produce a slice-weight matrix $\bm{W}_{N}=\{\bm{w}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times M}$, where $M$ is a hyperparameter denoting the number of slices. A channel-wise linear layer is subsequently applied to $\bm{H}_{N}$ to obtain $\bm{V}_{N}=\{\bm{v}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{h}}$. Using this slice-weight matrix, Physics-Attention computes a weighted average of $\bm{V}_{N}$ to obtain a slice feature matrix $\bm{S}_{M}=\{\bm{s}_{j}\}_{j=1}^{M}\in\mathbb{R}^{M\times d_{h}}$. This operation can be interpreted as projecting the original $N$ grid points onto $M$ physical slices, effectively compressing the sequence length from $N$ to $M$. The computation is formally given by:
@@
The $\text{Linear}^{(Q/K)}(\cdot)$ and $\text{Linear}^{(V)}(\cdot)$ are single-layer MLPs with output dimensions of $M$ and $d_{h}$, respectively. A self-attention on the slices $\bm{S}_{M}$ is then performed to produce an updated slices $\bm{S^{\prime}}_{M}=\{\bm{s^{\prime}}_{j}\}_{j=1}^{M}\in\mathbb{R}^{M\times d_{h}}$. Finally, using the slice-weight matrix $\bm{W}_{N}$, the slices $\bm{S^{\prime}}_{M}$ are decoded back to the original sequence length, yielding the output $\bm{H^{\prime}}_{N}=\{\bm{h^{\prime}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{h}}$. The process is summarized as follows:
@@
As shown in Eq. (6), linear attention aims to approximate the softmax function by designing different kernel functions $\varphi(\cdot)$ and $\psi(\cdot)$, allowing it to change the computation order of $\bm{Q}$, $\bm{K}$, and $\bm{V}$. In OFormer (2023a) and FactFormer (2023b), the kernel functions correspond to vector normalization and the identity function, respectively. The choice of kernel functions directly affects the performance of linear attention.
@@
Some improvements of linear attention insert additional intermediate operations between $\psi^{\top}(\bm{K})\bm{V}$ and $\varphi(\bm{Q})$, i.e., $\varphi(\bm{Q})\circ\mathcal{G}\circ\psi^{\top}(\bm{K})\bm{V}$. The operation $\mathcal{G}$ can be rank augmentation (Fan et al. 2025), forget gate (Yang et al. 2023), etc.
@@
In this section, we present a more generalized analysis of the Physics-Attention. Based on this, we redesign it through two key steps—generalization and simplification—to derive a more flexible architecture called the Linear Attention Neural Operator, or LinearNO. Figure 2(b) shows a comparison between the proposed LinearNO and the Physics-Attention.
@@
Luo et al. (2025) observed that Physics-Attention tends to produce uniform slice weights, thus generating less distinguishable slices in some cases, which damages the model’s performance, as shown in Figure 3. We hypothesize that this may be attributed to the intended asymmetry between $\varphi(\bm{Q})$ and $\psi(\bm{K})$ in linear attention; enforcing the shared learnable layer between $\varphi(\bm{Q})$ and $\psi(\bm{K})$ could blur their roles and result in overly averaged representations. Based on this hypothesis, we relaxed the weight-sharing constraint in the generalization step, resulting in asymmetric $\varphi(\bm{Q})$ and $\psi(\bm{K})$. Specifically, we allow $\psi(\bm{K})$ to be learned independently of $\varphi(\bm{Q})$. Following Physics-Attention, $\varphi(\bm{Q})$ and $\psi(\bm{K})$ are normalized along the $M$ and $N$ dimensions, respectively. This ensures that each row of attention matrix $\varphi(\bm{Q})\psi^{\top}(\bm{K})$ is normalized, which is consistent with full attention (Shen et al. 2021). Formally, we define $\varphi(\bm{Q})$ and $\psi(\bm{K})$ as follows:
@@
FNO (2021) / / / 0.1556 0.0108 /
U-FNO (2022) 0.0269 0.0056 0.0039 0.2231 0.0183 0.0239
GEO-FNO (2023c) 0.0138 0.0067 0.0074 0.1556 0.0108 0.0229
F-FNO (2023) 0.0078 0.0070 0.0047 0.2322 0.0077 0.0263
Galerkin (2021) 0.0118 0.0098 0.0120 0.1401 0.0084 0.0240
OFormer (2023a) 0.0183 0.0168 0.0017 0.1705 0.0124 0.0183
GNOT (2023) 0.0076 0.0047 0.0336 0.1380 0.0105 0.0086
FactFormer (2023b) 0.0071 0.0060 0.0312 0.1214 0.0109 /
ONO (2024) 0.0061 0.0052 0.0048 0.1195 0.0076 0.0118
LSM (2023) 0.0059 0.0050 0.0025 0.1535 0.0065 0.0218
LNO (2024) 0.0053 0.0031 0.0028 0.0830 0.0063 0.0066
Transolver (2024) 0.0053 0.0030 0.0013 0.0882 0.0055 0.0065
Transolver++ (2025) 0.0051 0.0027 0.0014 0.1010 0.0056 0.0064
@@
GraphSage (2017) 0.0087 0.0184 0.1476 0.9964 0.0461 0.1050 0.0270 0.9695
PointNet (2017) 0.0253 0.0996 0.1973 0.9919 0.0494 0.1104 0.0298 0.9583
GraphUNet (2019) 0.0076 0.0144 0.1677 0.9949 0.0471 0.1102 0.0226 0.9725
MeshGraphNet (2020) 0.0214 0.0387 0.2252 0.9945 0.0354 0.0781 0.0168 0.9840
GNO (2020) 0.0269 0.0405 0.2016 0.9938 0.0383 0.0815 0.0172 0.9834
GEO-FNO (2023c) 0.0361 0.0301 0.6161 0.9257 0.1670 0.2378 0.0664 0.8280
GALERKIN (2021) 0.0074 0.0159 0.2336 0.9951 0.0339 0.0878 0.0179 0.9764
GNOT (2023) 0.0049 0.0152 0.1992 0.9942 0.0329 0.0798 0.0178 0.9833
GINO (2023d) 0.0297 0.0482 0.1821 0.9958 0.0386 0.0810 0.0184 0.9826
LNO (2024) 0.0214 0.0268 0.1480 0.9744 0.0269 0.0870 0.0174 0.9781
Transolver (2024) 0.0023 0.0085 0.1230 0.9978 0.0221 0.0785 0.0117 0.9914
Transolver++ (2025) 0.0068 0.0159 0.1880 0.9910 0.0226 0.0800 0.0132 0.9914
@@
Datasets We adopt six benchmark datasets, including Airfoil, Pipe, Plasticity, NS, Darcy and Elasticity, which are classical problems in fluid mechanics and solid mechanics. These datasets were introduced by FNO (Li et al. 2021) and Geo-FNO (Li et al. 2023c), and are now widely used as standard benchmarks for evaluating Neural Operators. In addition, we use two industrial-level datasets: AirfRANS (Bonnet et al. 2022) and ShapeNet-Car (Umetani and Bickel 2018) to assess the model’s performance in aerodynamic shape design. More detailed information about all datasets can be found in Appendix B.
Baselines We compare our model against several representative baselines, including classical Neural Operator models: FNO (2021), Geo-FNO (2023c), F-FNO (2023) and U-FNO (2022), Transformer-based models: Galerkin (2021), OFormer (2023a), GNOT (2023), FactFormer (2023b), ONO (2024), LSM (2023), LNO (2024), Transolver (2024) and Transolver++ (2025), classical geometric deep models: PointNet (2017), GraphSage (2017) and GraphUNet (2019), MeshGraphNet (2020).
Setup To ensure fair comparisons, we follow the experimental settings of Transolver (Wu et al. 2024). We report the relative L2 error and Spearman correlation coefficient $\rho$ as the evaluation metric. Detailed experimental settings and hyperparameter configurations are provided in Appendix B.
Figure 4: All experiments are conducted on the Airfoil dataset. (a) and (b) show the final-layer slice-Weight matrix $W_{N}$ visualizations of LinearNO and Transolver, respectively. (c) shows the error distributions of LinearNO (top) and Transolver (middle), with the ground truth shown at the bottom. (d) shows the average rank of the attention matrix $\varphi(\bm{Q})\psi^{\top}(\bm{K})$ per head at each layer when slice numbers $M=64$. (e) presents the prediction errors of both models under different slice numbers.
@@
Table 1 presents the relative L2 errors of LinearNO and SOTA models across six PDE benchmarks. LinearNO achieves the best performance on all tasks. Specifically, LinearNO achieves over a 10% relative improvement on the Pipe, Plasticity, NS and Elasticity task.
@@
We compare the number of parameters and computational cost between LinearNO and Transolver on six standard PDE tasks, as shown in Table 3. It can be observed that LinearNO has significantly fewer parameters and lower computational cost than Transolver. On average, our method reduces the number of parameters by 40.0% and the FLOPs by 36.2%. This demonstrates that LinearNO is more computationally efficient and lightweight, making it more suitable for resource-constrained environments.
@@
Slice Analysis.
The rank of attention matrix $\varphi(\bm{Q})\psi^{\top}(\bm{K})$ corresponds to the number of physical states effectively captured by the model, which influences its performance (Luo et al. 2025). This effective rank is bounded by the number of slices $M$. We hypothesize that the symmetric design of $\varphi(\bm{Q})$ and $\psi(\bm{K})$ in Transolver limits the effective utilization of slices. In this section, we present a quantitative analysis to support this claim.
We estimate the layer-wise rank of the attention matrix using Singular Value Decomposition (Golub and Kahan 1965), and the results are shown in Figure 4. The findings indicate that the asymmetric design of $\varphi(\bm{Q})$ and $\psi(\bm{K})$ facilitates more efficient utilization of slices. Furthermore, we visualize the distribution of the final-layer $\varphi(\bm{Q})$ matrix on the Airfoil dataset. Compared to Transolver, LinearNO exhibits more diverse physical slices. In addition, LinearNO consistently outperforms Transolver across different numbers of slices on the Airfoil dataset.
@@
We design ablation experiments to evaluate the impact of the generalization and simplification steps on model performance. Specifically, we set $\psi^{\top}(\bm{K})=\text{Softmax}(\varphi^{\top}(\bm{Q}))$ to disable the generalization step, and we add a slice attention to disable the simplification step. We compare these variants with LinearNO, as shown in Table 4. By comparing No.1 vs. No.2 and No.3 vs. No.4, we observe that the generalization step consistently improves model performance in most cases, confirming the benefit of using asymmetric projections. Furthermore, the comparison between No.1 and No.3 shows that, under symmetric settings, slice attention does not consistently improve performance, which is also observed in Transolver. In contrast, comparing No.2 and No.4 reveals that applying slice attention under asymmetric projections consistently hurts the performance. Therefore, the simplification step is also justified.